• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 16:22:41 on September25,2005

    Entry number 155914

    keyword=Day Shift Summary

    Shift Summary: Day Shift Sept 25 2005

    Shift Leader/Target:   D.S. Armstrong
    3rd:   G. Urcioli
       At start of shift: continuing with production running, 38.2 µA on 4He target. Run 5354 continuing. Hall C is off completely (as is Hall B), so we are sole user at the moment. IHWP is IN. Right septum upper coil gets up to 7.70 K with 40 minutes of steady beam. Beam is very nice and steady. 9:45 Page Ryan, at the RC's suggestion, to arrange the remaining RHWP scans to take advantage of Hall C being off. Ryan explains to me what is needed, and we decide that Guido and I can do the scans ourselves, without requiring the experts (Ryan or Lisa) to come in... 10:35 We end Run 5357 early after 25 K good pairs, to start the source scan. RHWP Source Scan:
    1. end run 5357
    2. feedback off   (type this at command prompt on an apar@adaql account)
    3. dithering off   (using Green Monster)
    4. set Pockels cell voltages by hand to 180 V offsets using the "Laser Polarization & Parity Controls" medm screen:
       
      GUN3 PITA POS  voltage was 7.262 V, set to 4.262 V
       "     "  NEG  voltage was 4.063 V, set to 7.262 V
      
    5. 10:44 Start HAPPEX run 5358, IWHP in, RHWP scan
    6. once CODA started, invoke "auto_rhwp" at the ROC23 VxWorks window
    7. script starts, watch the RHWP update every 50 seconds in the Runbird window
    8. 11:06 script ends, end run 5358
    9. call MCC, ask to remove IHWP
    10. run flipper, but do not increment slug number (we are in slug 117)
    11. set Pockels cell voltages by hand:
       
      GUN3 PITA POS  voltage was 7.160 V, set to 4.160 V
       "     "  NEG  voltage was 4.400 V, set to 7.400 V
      
    12. 11:10 Start HAPPEX run 5359, IHWP OUT, RHWP scan
    13. once CODA started, invoke "auto_rhwp" at the ROC23 VxWorks window
    14. script starts, watch the RHWP update every 50 seconds in the Runbird window
    15. 11:32 script ends, end run.
    16. call MCC, insert IWHP again.
    17. run flipper, don't increment slug number.
    18. feedback on
    19. dithering (aka beam modulation) on
    20. 11:36 Start HAPPEX run 5360, production again
    Beam is marvelously stable throughout the scan.
    12:34 removed IHWP, ran flipper.
    Meanwhile, find that Compton was off, DAQ hangup; did a CODA reboot and restart. DAQ comes back up OK, and new run starts OK, but cavity power is zero. Laser is on, but cavity is not locked. Try the cavity lock instructions: Servo ON, slow Ramp procedure. No joy. Try a few more times, adjusting the "tracking level" down to 0.3 as per instructions. Still no joy. Page Sirish, but no answer yet. 13:30 Sirish calls; he has relocked cavity from home, but only at low power (150 W) instead of the 400 W we have had recently. Perhaps some laser resteering needs to be done? He will continue to work on it. At this low power our signal/noise would be rather marginal. 14:15 Sirish has resteered the Compton laser beam using the mirrors, and now the cavity power is around 500 W. I start a new Compton run (12297).
    15:45 Did a scanner run (run 5365).
    16:00 Hand off to next shift, production run 5365 is underway; some trouble with that silly piece of software known as BTA.

    HAPPEX Runs Taken

       Run     Pairs    IHWP
    
       5354     45 K     IN   Production
       5355     48 K     IN   Production
       5356     48 K     IN   Production
       5357     25 K     IN   Production
       5358     18 K     IN   RWHP source scan - not production
       5359     20 K     OUT  RWHP source scan - not production
       5360     49 K     IN   Production
       5361     48 K     OUT  Production - start of slug 118
       5362     46 K     OUT  Production 
       5363     33 K     OUT  Production 
       5364              OUT  Q2 scanner run - not production      
       5365  (underway)  OUT  Production   
    

    Compton Runs Taken

       Run 12297   82.2% polarization, signal/background = 7.2
       Run 12298      (still underway)